calculate the permutation 60P47. enter you answer using E-notation where the mantissa is between 1 and 10 and rounded to two decimal places
Answer #1

Solution

We have to calculate permutation _{60} P_{47}

Formula to solve permutation as follow

_{n} P_{r}= \frac{n!}{(n-r)!}

So, _{60} P_{47}= \frac{60!}{(60-47)!}= 1.133627*10^{72}

So, _{60} P_{47}= = 1.13E+10^{72}